How do i turn this into a y= equation?
x+3y=-6 and x=3y

Answer by checkley75(3666)

x+3y=-6
3y=-x-6
y=-x/3-6/3
y=-x/3-2 answer.
----------------------
x=3y
y=x/3 answer.